# Age-Optimal Sampling and Routing under Intermittent Links and Energy Constraints

### v1

@@ III-A Fractional programming reformulation
$\limsup_{n\to\infty}\Big(\mathbb{E}\left[\frac{1}{n}\textstyle\sum_{i=0}^{n-1}(E_{i+1}-E_{\max}Z_{i}-E_{\max}Y_{i+1,R_{i}})\right]\leq 0.$ (11)
@@
$\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}\!\Big[q(Y_{i,R_{i-1}},Y_{i+1,R_{i}},Z_{i})-\lambda(Z_{i}+Y_{i+1,R_{i}})\Big]\bigg\}$ (12a)
$\text{s.t.}\limsup_{n\to\infty}\Big(\mathbb{E}\left[\frac{1}{n}\textstyle\sum_{i=0}^{n-1}(E_{i+1}-E_{\max}(Z_{i}+Y_{i+1,R_{i}}))\right]\leq 0$ (12b)
The function $h(\lambda)$ measures the difference between the expected accumulated AoI cost and the weighted accumulated epoch duration. The problem is a CMDP, we next transform it into a standard MDP utilizing the Lagrangian techniques.
@@ III-B Lagrangian Techniques
$H(\lambda,c)\triangleq\min_{\pi\in\Pi}\bigg\{\limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}\!\Big[q(Y_{i,R_{i-1}},Y_{i+1,R_{i}},Z_{i})$ (13)
$-\lambda(Z_{i}+Y_{i+1,R_{i}})-c\big(E_{\max}(Z_{i}+Y_{i+1,R_{i}})-E_{i+1}\big)\Big]\bigg\},$
@@
The above formulation can be interpreted as the Lagrangian relaxation of the constrained problem.
For a fixed $c$, solving $h(\lambda,c)=0$ yields the conditionally optimal long-term average AoI $\lambda_{c}^{\star}$. Let $E(\lambda_{c}^{\star},c)$ denote the corresponding average energy consumption of the conditionally optimal policy which we denote by $\pi_{c}^{\star}$. Then, the average energy consumption is:
$E(\lambda_{c}^{\star},c)\triangleq\limsup_{n\to\infty}\frac{\displaystyle\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}\big(G_{r_{i,c}^{\star}}\,Y_{i+1,R_{i}}+C_{s}\big)\right]}{\displaystyle\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(z_{i,c}^{\star}+Y_{i+1,R_{i}})\right]},$ (14)
where $r_{i,c}^{\star}$ and $z_{i,c}^{\star}$ are $i$-th epoch routing and waiting decisions made by using $\pi_{c}^{\star}$. Since $E(\lambda_{c}^{\star},c)$ is monotonically decreasing in $c$ by complementary slackness [altman2021constrained], the optimal Lagrange multiplier is obtained as
$c^{\star}=\inf\{\,c>0:E(\lambda_{c}^{\star},c)\leq E_{\max}\,\}.$ (15)
At $(\lambda^{\star},c^{\star})$, both the average AoI and energy constraint are simultaneously optimal, satisfying $h(\lambda^{\star},c^{\star})=0$.
@@
See Appendix A. ∎
@@ III-C Average-Cost MDP for a Given $\lambda$ and $c$
In the third step, we show that Problem 3 can be formulated as an average-cost Markov Decision Process (MDP) for a fixed value of $\lambda$, described by the quadruple
$\mathscr{M}(\lambda)\triangleq(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{C}).$ (16)
Each component of this MDP is defined as follows.
  • State Space $\mathcal{S}=[0,\infty)\times\{0,1\}^{N}$: At each decision epoch $i$, the system occupies a state
    $(Y_{i,R_{i-1}},\mathbf{L}_{i})\in\mathcal{S},$ (17)
    where the continuous component $Y_{i,R_{i-1}}=y<\infty$ denotes the delay observed in the previous transmission, and
    $\mathbf{L}_{i}=(l_{1,i},l_{2,i},\ldots,l_{N,i})\in\{0,1\}^{N}$ (18)
    encodes the availability of the $N$ routes at epoch $i$. Specifically,
@@
    The set of admissible actions depends on $\mathbf{L}_{i}$:
    $\mathcal{A}(\mathbf{L}_{i})=\mathcal{R}_{i}\times\mathbb{R}^{+}$ (20)
@@
    The global action space is the union of all state-dependent sets:
    $\mathcal{A}=\bigcup_{\mathbf{L}\in\{0,1\}^{N}}\mathcal{A}(\mathbf{L}).$ (21)
@@
  • State Transition Probability $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathscr{B}(\mathcal{S})\to[0,1]$: Let $\mathscr{B}(\mathcal{S})$ denote the Borel $\sigma$-algebra generated by the measurable subsets of $\mathcal{S}=[0,\infty)\times\{0,1\}^{N}$. For any current state $(y,\mathbf{l})\in\mathcal{S}$, admissible action $(r,z)\in\mathcal{A}(\mathbf{l})$, and measurable set $C\in\mathscr{B}(\mathcal{S})$, the transition kernel is defined as $\mathcal{P}\big(C\mid y,\mathbf{l},r,z\big)=\mathbb{P}\big((Y_{i+1,R_{i}},\mathbf{L}_{i+1})\in C\mid Y_{i,R_{i-1}}=y,\mathbf{L}_{i}=\mathbf{l},R_{i}=r,Z_{i}=z\big).$ We assume that, conditioned on the current state and action, the next delay $Y_{i+1}$ and the next availability vector $\mathbf{L}_{i+1}$ are conditionally independent. That is, $\mathbb{P}\big(Y_{i+1,R_{i}},\mathbf{L}_{i+1}\mid Y_{i,R_{i-1}}=y,\mathbf{L}_{i}=\mathbf{l},R_{i}=r,Z_{i}=z\big)=\mathbb{P}(Y_{i+1.R_{i}}\mid R_{i}=r)\mathbb{P}(\mathbf{L}_{i+1}),$ meaning that the delay distribution depends only on the chosen route $r$, while the availability process evolves independently of the previous state or action. Hence, for any measurable set $C\subseteq\mathcal{S}$,
    $\mathcal{P}\big(C\mid y,\mathbf{l},r,z\big)=$ (22)
    $\sum_{\mathbf{l}^{\prime}\in\{0,1\}^{N}}\mathbb{P}(\mathbf{L}_{i+1}=\mathbf{l}^{\prime})\int_{C_{Y}(\mathbf{l}^{\prime})}Q_{r}(y^{\prime})dy^{\prime},$
    where $C_{Y}(\mathbf{l}^{\prime}):={y^{\prime}\geq 0:(y^{\prime},\mathbf{l}^{\prime})\in C}.$ The availability process $\mathbf{L}_{i}$} is independent across epochs and identically distributed with stationary distribution
    $\mathbb{P}(\mathbf{L}_{i+1}=\mathbf{l}^{\prime})=\prod_{k=1}^{N}p_{k}^{1-l_{k}^{\prime}}(1-p_{k})^{l_{k}^{\prime}}.$ (23)
@@
  • Cost Function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$: The one-step cost incurred when the system is in state $(y,\mathbf{l})$ and action $(r,z)$ is taken is denoted by $g(y,\mathbf{l},z,r;\lambda,c)$, defined as
    $g(y,\mathbf{l},z,r;\lambda,c)=\mathbb{E}_{Y\sim Q_{r}}\!\left[\frac{(2y+Y+z)(Y+z)}{2}\right]+cC_{s}$ (24)
    $-(\lambda+cE_{\max})z-\left(\lambda+cE_{\max}-cG_{r}\right)\,\mathbb{E}_{Y\sim Q_{r}}[Y].$
    By substituting $\mu_{r}=\mathbb{E}_{Y\sim Q_{r}}[Y]$ and $\sigma_{r}^{2}=\mathrm{Var}_{Y\sim Q_{r}}[Y]$, this simplifies to
    $g(y,\mathbf{l},z,r;\lambda,c)=\frac{z^{2}}{2}+(y+\mu_{r}-\lambda-cE_{\max})\,z$ (25)
    $+(y+G_{r}-cE_{\max}-\lambda)\mu_{r}+\frac{\mu_{r}^{2}+\sigma_{r}^{2}}{2}+cC_{s}.$
@@
Using Lemma 1 and the constructed MDP $\mathscr{M}(\lambda)$ for a fixed $\lambda$, we can design a nested three-layer optimization algorithm (e.g., [bedewy2021optimal, 10806969]) to solve the CMDP in Problem (1). The details of this numerical solution are presented in Section V.
@@ IV-A Structural Results of Optimal Policies
The following Theorem 1 establishes the piecewise-threshold structure of the optimal sampling and routing policies.
@@
For an $N$-route problem where the mean delay of each route satisfies $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{N}$ and the delay distribution of each route has infinite support, the jointly optimal sampling and routing policies exhibit the following threshold structure:
@@
    Optimal Routing: The optimal routing action at the $i$-th epoch, denoted by $R_{i}^{\star}$, is a monotonic non-decreasing step function of the observed delay $Y_{i,R_{i-1}}$, and can be determined by $K\leq|\mathcal{R}_{i}|-1$ positive thresholds $0<\tau_{1}(\mathbf{l}_{i})<\tau_{2}(\mathbf{l}_{i})\cdots<\tau_{K}(\mathbf{l}_{i})$ and $K+1$ monotonic increasing index values $a_{1}(\mathbf{l}_{i})<a_{2}(\mathbf{l}_{i})<\cdots<a_{K+1}(\mathbf{l}_{i})\in\mathcal{R}_{i}$:
    $R_{i}^{\star}(\mathbf{l}_{i})=\sum_{k=1}^{K+1}\big(a_{k}(\mathbf{l}_{i})-a_{k-1}(\mathbf{l}_{i})\big)u\big(Y_{i,R_{i-1}}-\tau_{k-1}(\mathbf{l}_{i})\big),$ (26)
    where $\tau_{0}(\mathbf{l}_{i})\triangleq 0$, $a_{0}(\mathbf{l}_{i})\triangleq 0$, and $u(t)$ is the unit step function:
    $u(t)\triangleq\begin{cases}0,&t<0\\ 1,&t\geq 0.\end{cases}$ (27)
    Furthermore, the number of unique thresholds is upper bounded by $\frac{N(N-1)}{2}$.
@@
    Optimal Sampling: The optimal waiting time at the $i$-th epoch, denoted by $Z_{i}^{\star}$, follows a water-filling structure and can be determined by $K+1$ thresholds $\beta_{1}^{\star}(\mathbf{l}_{i})<\beta_{2}^{\star}(\mathbf{l}_{i})<\cdots<\beta_{K+1}^{\star}(\mathbf{l}_{i})$ with $\beta_{k}^{\star}(\mathbf{l}_{i})=\lambda^{\star}-\mu_{a_{k}(\mathbf{l}_{i})}$,
    $Z_{i}^{\star}(\mathbf{l}_{i})=$ (28)
[…]
    or equivalently,
    $Z_{i}^{\star}(\mathbf{l}_{i})=(\lambda^{\star}+c^{\star}E_{\max}-\mu_{R_{i}^{\star}(\mathbf{l}_{i})}-Y_{i,R_{i-1}})^{+},$ (29)
    where $\lambda^{\star}$ is the optimal average AoI defined in Problem 1, and $(\cdot)^{+}$ is defined as $(\cdot)^{+}\triangleq\max\{0,\cdot\}$.
@@
With the MDP $\mathscr{M}(\lambda)$ (25), we can establish the Average-Cost Optimality Equation (ACOE) [howard1960dynamic, Eq. 4.1]:
@@
where $V^{*}(y,\mathbf{l};\lambda,c)$ is the relative value function, and $h(\lambda,c)$ is the optimal value of the reformulated MDP in Problem 3. Given any $\lambda$, $c$ and route $r\in\mathcal{N}$, we first prove that the optimal waiting time that solves the right hand-side of (30) follows a water-filling structure, given by:
$z^{\star}(y,\mathbf{l};r,\lambda,c)=(\lambda+cE_{\max}-\mu_{r}-y)^{+}.$ (31)
@@
Then, we analyze a series of properties of the function $Q(y,\mathbf{l},r)$ and prove that $r^{\star}(y,\mathbf{l})$ is a non-decreasing step function, thus accomplishing the proof. See Section VI for the detailed proof. ∎
We have established that the AoI-optimal routing policy follows a threshold-based structure. However, one can argue that these thresholds never actually exist and that the optimal policy always uses a single route. To counter this, we show that there exist system configurations where these thresholds must exist.
The following Lemma 2 demonstrates an important relationship between the sampling threshold $\beta_{k}^{\star}(\mathbf{l}_{i})$ and the routing threshold $\tau_{k}(\mathbf{l}_{i})$.
@@
The following assertion holds true:
@@
Consequently, for any interval $Y_{i,R_{i-1}}\in\big[\tau_{k-1}(\mathbf{l}_{i}),\tau_{k}(\mathbf{l}_{i})\big)$ associated with a constant optimal routing option $a_{k}(\mathbf{l}_{i})$, there exists a corresponding sub-interval $\big[\beta^{\star}_{k}(\mathbf{l}_{i}),\tau_{k}(\mathbf{l}_{i})\big)$ in which a zero-waiting policy, defined by $Z_{i}^{\star}(\mathbf{l}_{i})=\big(\beta_{k}^{\star}(\mathbf{l}_{i})-Y_{i,R_{i-1}}\big)^{+}=0$, is optimal. Figure 3 provides an example sketch that illustrates the structure of the jointly optimal sampling and routing policy under a given availability state $\mathbf{l}$. The threshold-based structure derived in this subsection enables highly efficient deployment in complex networks. Terminals can maximize information freshness simply by storing and applying the derived thresholds. In Section V, we present a series of algorithms to compute these thresholds efficiently.
@@
IV-B Bounds on the Optimal Average AoI
In this subsection, we establish the upper and lower bounds on the optimal average age $\lambda^{\star}$. These bounds will later serve as initialization points for the bisection search described in Section V.

### v2

@@ III-A Fractional programming reformulation
$\limsup_{n\to\infty}\Big(\mathbb{E}\left[\frac{1}{n}\textstyle\sum_{i=0}^{n-1}(E_{i+1}-E_{\max}Z_{i}-E_{\max}Y_{i+1})\right]\leq 0.$ (9)
@@
$\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}\!\Big[q(Y_{i},Z_{,}i,Y_{i+1})-\lambda(Z_{i}+Y_{i+1})\Big]\bigg\}$ (10a)
$\text{s.t.}\limsup_{n\to\infty}\Big(\mathbb{E}\left[\frac{1}{n}\textstyle\sum_{i=0}^{n-1}(E_{i+1}-E_{\max}(Z_{i}+Y_{i+1}))\right]\leq 0$ (10b)
The function $h(\lambda)$ measures the difference between the expected accumulated AoI cost and the weighted accumulated epoch duration. This problem is formulated as a CMDP, which we next transform into a standard MDP by utilizing Lagrangian techniques.
@@ III-B Lagrangian Techniques
$h(\lambda,c)\triangleq\min_{\pi\in\Pi}\bigg\{\limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}\!\Big[q(Y_{i},Z_{i},Y_{i+1})$ (11)
$-\lambda(Z_{i}+Y_{i+1})-c\big(E_{\max}(Z_{i}+Y_{i+1})-E_{i+1}\big)\Big]\bigg\},$
@@
The proof closely follows the derivation in [30, Lemma 2], adapted here for the constrained Lagrangian formulation. See Appendix A. ∎
Hence, for a fixed $c$, solving $h(\lambda,c)=0$ yields the conditionally optimal long-term average AoI $\lambda_{c}^{\star}$. Let $E(\lambda_{c}^{\star},c)$ denote the corresponding average energy consumption of the conditionally optimal policy which we denote by $\pi_{c}^{\star}$. Then, the average energy consumption is:
$E(\lambda_{c}^{\star},c)\triangleq\limsup_{n\to\infty}\frac{\displaystyle\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}\big(G_{r_{i,c}^{\star}}\,Y_{i+1}+C_{s}\big)\right]}{\displaystyle\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(z_{i,c}^{\star}+Y_{i+1})\right]},$ (12)
where $r_{i,c}^{\star}$ and $z_{i,c}^{\star}$ are $i$-th epoch routing and waiting decisions made by using $\pi_{c}^{\star}$. Since $E(\lambda_{c}^{\star},c)$ is monotonically decreasing in $c$ by complementary slackness [1], the optimal Lagrange multiplier is obtained as $c^{\star}=\inf\{\,c>0:E(\lambda_{c}^{\star},c)\leq E_{\max}\,\}.$ At $(\lambda^{\star},c^{\star})$, both the average AoI and energy constraint are simultaneously optimal, satisfying $h(\lambda^{\star},c^{\star})=0$.
@@ III-C Average-Cost MDP for a Given $\lambda$ and $c$
In the third step, we show that Problem 3 can be formulated as an average-cost Markov Decision Process (MDP) for a fixed value of $\lambda$, described by the quadruple $\mathscr{M}(\lambda)\triangleq(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{C}).$ Each component of this MDP is defined as follows.
  • State Space $\mathcal{S}=[0,\infty)\times\{0,1\}^{N}$: At each decision epoch $i$, the system occupies a state $(Y_{i},\mathbf{L}_{i})\in\mathcal{S}$, where the continuous component $Y_{i}=y<\infty$ denotes the delay observed in the previous transmission, and $\mathbf{L}_{i}=(l_{1,i},l_{2,i},\ldots,l_{N,i})\in\{0,1\}^{N}$ encodes the availability of the $N$ routes at epoch $i$. Specifically,
@@
    The set of admissible actions depends on $\mathbf{L}_{i}$: $\mathcal{A}(\mathbf{L}_{i})=\mathcal{R}_{i}\times\mathbb{R}^{+}$. The global action space is defined as the union of all state-dependent sets: $\mathcal{A}\triangleq\bigcup_{\mathbf{L}\in\{0,1\}^{N}}\mathcal{A}(\mathbf{L})$.
@@
  • State Transition Probability $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathscr{B}(\mathcal{S})\to[0,1]$: Let $\mathscr{B}(\mathcal{S})$ denote the Borel $\sigma$-algebra generated by the measurable subsets of $\mathcal{S}=[0,\infty)\times\{0,1\}^{N}$. For any current state $(y,\mathbf{l})\in\mathcal{S}$, admissible action $(r,z)\in\mathcal{A}(\mathbf{l})$, and measurable set $C\in\mathscr{B}(\mathcal{S})$, the transition kernel is defined as $\mathcal{P}\big(C\mid y,\mathbf{l},r,z\big)=\mathbb{P}\big((Y_{i+1},\mathbf{L}_{i+1})\in C\mid Y_{i}=y,\mathbf{L}_{i}=\mathbf{l},R_{i}=r,Z_{i}=z\big).$ We assume that, conditioned on the current state and action, the next delay $Y_{i+1}$ and the next availability vector $\mathbf{L}_{i+1}$ are conditionally independent. That is, $\mathbb{P}\big(Y_{i+1},\mathbf{L}_{i+1}\mid Y_{i}=y,\mathbf{L}_{i}=\mathbf{l},R_{i}=r,Z_{i}=z\big)=\mathbb{P}(Y_{i+1}\mid R_{i}=r)\mathbb{P}(\mathbf{L}_{i+1}),$ meaning that the delay distribution depends only on the chosen route $r$, while the availability process evolves independently of the previous state or action. Hence, for any measurable set $C\subseteq\mathcal{S}$, $\mathcal{P}\big(C\mid y,\mathbf{l},r,z\big)=\sum_{\mathbf{l}^{\prime}\in\{0,1\}^{N}}\mathbb{P}(\mathbf{L}_{i+1}=\mathbf{l}^{\prime})\int_{C_{Y}(\mathbf{l}^{\prime})}Q_{r}(y^{\prime})dy^{\prime}$, where $C_{Y}(\mathbf{l}^{\prime}):={y^{\prime}\geq 0:(y^{\prime},\mathbf{l}^{\prime})\in C}.$ The availability process $\mathbf{L}_{i}$} is independent across epochs and identically distributed with stationary distribution $\mathbb{P}(\mathbf{L}_{i+1}=\mathbf{l}^{\prime})=\prod_{k=1}^{N}p_{k}^{1-l_{k}^{\prime}}(1-p_{k})^{l_{k}^{\prime}}$.
@@
  • Cost Function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$: Let $F(t)\triangleq\int_{0}^{t}f(\tau)d\tau$ be the primitive of the penalty function $f(\cdot)$. The one-step cost incurred when the system is in state $(y,\mathbf{l})$ and action $(r,z)$ is taken is denoted by $g(y,\mathbf{l},z,r;\lambda,c)$, defined as
    $g(y,\mathbf{l},z,r;\lambda,c)$ (14)
    $\triangleq\mathbb{E}_{Y\sim Q_{r}}\!\left[q(y,z,Y)\right]+cC_{s}-(\lambda+cE_{\max})z$
    $-\left(\lambda+cE_{\max}-cG_{r}\right)\,\mathbb{E}_{Y\sim Q_{r}}[Y]$
    $=\mathbb{E}_{Y\sim Q_{r}}\!\left[F(y+Y+z)-F(y)\right]+cC_{s}-(\lambda+cE_{\max})z$
    $-\left(\lambda+cE_{\max}-cG_{r}\right)\,\mu_{r}.$
    Here, the term $q(y,z,Y)$ represents the accumulated penalty area during the updating interval.
@@
Using Lemma 1 and the constructed MDP $\mathscr{M}(\lambda)$ for a fixed $\lambda$, we can design a nested three-layer optimization algorithm (e.g., [5, 20]) to solve the CMDP in Problem (1). The details of this numerical solution are presented in Section V.
@@ IV-A Structural Results of Optimal Policies
The introduction of general non-linear penalty functions $f(\cdot)$ generalizes the optimal policy structure. While the optimal waiting policy maintains a target-threshold structure for any monotonic non-decreasing $f(\cdot)$, the strict monotonicity of the routing policy is guaranteed only under certain conditions.
@@
For an $N$-route problem where the delay distribution of each route has infinite support, the jointly optimal sampling and routing policies exhibit the following structure:
@@
    Optimal Routing (Conditional): Let $\Psi_{r}(y)\triangleq\mathbb{E}_{Y\sim Q_{r}}[f(y+Y)]$ denote the expected immediate penalty on route $r$ given a current system delay $y$. If the penalty function $f(\cdot)$ and the delay distributions satisfy the Expected Penalty Ordering property, defined as:
    $\mu_{j}\geq\mu_{k}\implies\Psi_{j}(y)\geq\Psi_{k}(y)\quad\forall y\geq 0,$ (15)
    then the optimal routing action $R_{i}^{\star}$ is a monotonic step function of the observed delay $Y_{i}$. Specifically, the routing decision is determined by $K$ positive thresholds such that the policy switches to routes with statistically ”better” (lower mean) delay profiles as the previous delay $Y_{i}$ increases:
    $R_{i}^{\star}(\mathbf{l}_{i})=\sum_{k=1}^{K+1}\big(a_{k}(\mathbf{l}_{i})-a_{k-1}(\mathbf{l}_{i})\big)u\big(Y_{i}-\tau_{k-1}(\mathbf{l}_{i})\big),$ (16)
    where $a_{k}(\mathbf{l}_{i})$ are indices of routes sorted by preference, and $u(t)$ is the unit step function.
    Examples: This sufficient condition is satisfied by:
    • The Linear penalty $f(\Delta)=\Delta$ (since $\mu_{j}>\mu_{k}\implies y+\mu_{j}>y+\mu_{k}$).
    • The Exponential penalty $f(\Delta)=e^{\alpha\Delta}-1$, provided the delay distributions satisfy the monotonicity of the moment generating function (i.e., $\mu_{j}>\mu_{k}\implies M_{j}(\alpha)>M_{k}(\alpha)$).
    Remark 2.
    For general convex functions (e.g., $f(\Delta)=\Delta^{2}$), this monotonicity is not guaranteed due to the coupling of current age with higher-order delay moments.
@@
    Optimal Sampling (General): For any monotonic non-decreasing penalty function $f(\cdot)$, the optimal waiting time at the $i$-th epoch, $Z_{i}^{\star}$, follows a target-threshold structure. Specifically, for a chosen route $R_{i}$, there exists an optimal target age $\beta^{\star}(R_{i})$ such that:
    $Z_{i}^{\star}(\mathbf{l}_{i})=\left(\beta^{\star}(R_{i})-Y_{i}\right)^{+},$ (17)
    where $(\cdot)^{+}\triangleq\max\{0,\cdot\}$. The target age threshold $\beta^{\star}(R_{i})$ is the unique solution to the following stopping rule:
    $\beta^{\star}(R_{i})=$ (18)
    $\inf\left\{t\geq 0:\mathbb{E}_{Y}\left[f(t+Y_{i+1})\right]\geq\lambda^{\star}+c^{\star}E_{\max}\right\},$
    where $Y_{i+1}$ is the random delay of the chosen route, $\lambda^{\star}$ is the optimal average AoI penalty, and $c^{\star}$ is the Lagrange multiplier associated with the energy constraint.
    This implies the system waits until the expected penalty at the moment of delivery reaches the critical system cost $\lambda^{\star}+c^{\star}E_{\max}$. In the special case where $f(\Delta)=\Delta$ (linear), (18) simplifies to $\beta^{\star}(R_{i})=\lambda^{\star}+c^{\star}E_{\max}-\mu_{R_{i}}$, recovering the standard result.
@@
With the MDP $\mathscr{M}(\lambda)$ defined by the cost function in (14), we can establish the Average-Cost Optimality Equation (ACOE) [13, Eq. 4.1]:
@@
where $V^{*}(y,\mathbf{l};\lambda,c)$ is the relative value function, and $h(\lambda,c)$ is the optimal value of the reformulated MDP in Problem 3. Given any $\lambda$, $c$ and route $r\in\mathcal{N}$, we first prove that the optimal waiting time that solves the right hand-side of (19) follows a target-threshold structure, given by:
$z^{\star}(y,\mathbf{l};r,\lambda,c)=(\beta^{\star}(r,\lambda,c)-y)^{+},$ (20)
where $\beta^{\star}(r,\lambda,c)$ is the route-specific target age derived from the stopping rule given in (18).
@@
Then, we analyze a series of properties of the function $Q(y,\mathbf{l},r)$ and prove that, under the Expected Penalty Ordering property, $r^{\star}(y,\mathbf{l})$ is a non-decreasing step function. The full derivation of this result is provided in Section VI. ∎
The following Lemma 2 demonstrates an important relationship between the sampling target threshold $\beta_{k}^{\star}(\mathbf{l}_{i})$ (associated with the $k$-th preferred route $a_{k}$) and the routing switching threshold $\tau_{k}(\mathbf{l}_{i})$.
@@
For penalty functions yielding a monotonic routing structure, the following assertion holds true:
@@
This implies that the optimal policy always terminates the waiting phase before the age accumulates to the level where a route switch would become optimal.
@@
Consequently, for such penalty functions, for any interval $Y_{i}\in\big[\tau_{k-1}(\mathbf{l}_{i}),\tau_{k}(\mathbf{l}_{i})\big)$ associated with a constant optimal routing option $a_{k}(\mathbf{l}_{i})$, there exists a corresponding sub-interval $\big[\beta^{\star}_{k}(\mathbf{l}_{i}),\tau_{k}(\mathbf{l}_{i})\big)$ in which a zero-waiting policy, defined by $Z_{i}^{\star}(\mathbf{l}_{i})=\big(\beta_{k}^{\star}(\mathbf{l}_{i})-Y_{i}\big)^{+}=0$, is optimal. Figure 4 provides an example sketch that illustrates the structure of the jointly optimal sampling and routing policy under a given availability state $\mathbf{l}$. The threshold-based structure derived in this subsection enables highly efficient deployment in complex networks. Terminals can maximize information freshness simply by storing and applying the derived thresholds. In Section V, we present a series of algorithms to compute these thresholds efficiently.
@@
IV-B Bounds on the Optimal Average AoI Penalty
In this subsection, we establish the upper and lower bounds on the optimal average age penalty $\lambda^{\star}$. These bounds will later serve as initialization points for the bisection search described in Section V.